OpenStudy (anonymous):

6Xsquared-41x-7 how do i factor this??

OpenStudy (anonymous):

You can find its roots, say x1 and x2 and then write it in the form a(x-x1)(x-x2), where a is the coefficient of x^2 (6 in this case).

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

im somewhat confused so what would go in parenthesis, (6x-x)(x-x) what are the missing x's?

OpenStudy (anonymous):

The roots of the quadratic. You find the discriminant and after that, you find the roots. Do you know the methodology?

OpenStudy (anonymous):

no those words are too big

OpenStudy (anonymous):

A quadratic is written in the form: \[ax^2+bx+c\] Here you have: \[6x^2-41x-7\] So what would be the values for a,b,c in this case?

OpenStudy (anonymous):

\[6x ^{2}-41x-7=6x ^{2}-\left( 42-1 \right)x-7 =6x \left( x-7 \right)+1\left( x-7 \right)\] \[=\left( x-7 \right)\left( 6x+1 \right)\]